OpenStudy (anonymous):

Slope -2, passes through (-3, 14)

OpenStudy (shamil98):

y - y1 = m(x-x1) plug it in and solve.

OpenStudy (anonymous):

It just said put the equation

OpenStudy (anonymous):

Okay captainmico remember from you last question what is the m value?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

okay now can you write the equation subing in correct values for x,y, and m but leaving the b value alone?

OpenStudy (anonymous):

Y=-2x14-3?

OpenStudy (anonymous):

not quite lets look at the coordinates they gave you (-3,14) these coordinates mean that if x=-3 then y=14

OpenStudy (anonymous):

you already said that m=-2

OpenStudy (anonymous):

so sub in the known values for x,y and m into the general equation y=mx+b

OpenStudy (anonymous):

14=-2-3x

OpenStudy (anonymous):

Or the x is in 2

OpenStudy (anonymous):

...no look lets take this one step at a time lets look at the general equation again.y=mx+b this time sub in only the y value for the moment

OpenStudy (anonymous):

Okay so -2x - 3=14

OpenStudy (anonymous):

...look I will give you the substitution this time y=mx+b 14=-2*-3+b

OpenStudy (anonymous):

can you solve that equation for b?

OpenStudy (anonymous):

-5? How?

OpenStudy (anonymous):

...can you post the steps you took in solving that equation for me?

OpenStudy (anonymous):

20? I need to finish this asap

OpenStudy (anonymous):

very very close look you multiply -2 and -3 together to get +6 so you have 14=6+b

OpenStudy (anonymous):

what is your next step?

OpenStudy (anonymous):

Yeah i minused it

OpenStudy (anonymous):

so what is the value of b?

OpenStudy (anonymous):

Normally I wouldn't tell you the correct answer like this but...nobody else is helping you with this one and I really have to go so b=8